# Character theoretic techniques for nonabelian partial difference sets

### v1

@@ 4.1. Modular computation of conjugacy class intersection sizes for PDSs
[0,9,1,9,1,1,1,1,1,1,1,1,1,1,1][0,9,1,9,1,1,1,1,1,1,1,1,1,1,1][ 0 , 9 , 1 , 9 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 ]
as a potentially valid list of conjugacy class intersection sizes for a (111,30,5,9)(111,30,5,9)( 111 , 30 , 5 , 9 )-PDS in SmallGroup(111, 1). Note that a subset of this group with these intersection sizes will evaluate on irreducible characters to values admissible by Lemma 3.2. However, by turning the conjugacy class and PDS constraints into a linear program, we are able to rule out the existence of such a PDS using the linear optimization software Gurobi [19], and hence there does not exist a (111,30,5,9)(111,30,5,9)( 111 , 30 , 5 , 9 )-PDS.
@@
[0,4,14,1,1,4,4,1,2,4],[0,4,14,1,1,4,4,2,1,4][0,4,14,1,1,4,4,1,2,4],[0,4,14,1,1,4,4,2,1,4][ 0 , 4 , 14 , 1 , 1 , 4 , 4 , 1 , 2 , 4 ] , [ 0 , 4 , 14 , 1 , 1 , 4 , 4 , 2 , 1 , 4 ]
as potentially valid lists of conjugacy class intersections for a (96,35,10,14)(96,35,10,14)( 96 , 35 , 10 , 14 )-PDS in SmallGroup(96, 64) and SmallGroup(96, 227), respectively. These were also ruled out with the use of Gurobi [19], and, since these are the only valid lists of conjugacy class intersection sizes for these parameters, there does not exist a (96,35,10,14)(96,35,10,14)( 96 , 35 , 10 , 14 )-PDS, either. The authors would be quite interested in a proof of the nonexistence of these PDSs that does not use the brute force of linear programming.
@@ 4.3. PDSs constructed using Algorithm CCI
For a number of groups, we used this method to calculate all possible sizes of intersections of the conjugacy classes of the group with a possible PDS inside this group. We then ran a restricted search for PDSs satisfying the class intersection restrictions using a modified Random Restart Hill Climb search described in [2]. Using the technique from [2], we were able to construct nineteen distinct PDSs in nonabelian groups, many of which were already explicitly recognized, such as PDSs arising from triangular graphs. We found three parameter sets, (111,44,19,16)(111,44,19,16)( 111 , 44 , 19 , 16 ), (305,76,27,16)(305,76,27,16)( 305 , 76 , 27 , 16 ), and (981,140,43,16)(981,140,43,16)( 981 , 140 , 43 , 16 ), which correspond to an infinite family of groups acting regularly on Steiner systems ${\mathrm{S}}(2,4,p^{d})$ for $p$ prime. Five others that we were able to construct computationally – in SmallGroup(57, 1), SmallGroup(155, 1), SmallGroup(301, 1), SmallGroup(737, 1), and SmallGroup(1027, 1) – have parameters corresponding to an infinite family of PDSs corresponding to Steiner triple systems. The infinite families on ${\mathrm{S}}(2,4,p)$ and ${\mathrm{S}}(2,3,p^{d})$ correspond to block-regular Steiner 222-design constructions given by Clapham in [6] and Wilson in [33]. Finally, we found a (183,70,29,25)(183,70,29,25)( 183 , 70 , 29 , 25 )-PDS in SmallGroup(183, 1) which is related to a ${\mathrm{S}}(2,5,61)$ Steiner system and, as it turns out, is also part of an infinite family. To the authors knowledge, the PDSs arising from [33] have not previously been recognized in the literature, and we are only aware of one paper [28] recognizing the constructions given by Clapham [6] as giving rise to PDSs.
@@
Let $p$ be a prime such that $p^{d}>9$ and $p^{d}\equiv 7\pmod{12}$. Then, there exist $(p^{d}(p^{d}-1)/6,3(p^{d}-3)/2,(p^{d}+3)/2,9)$-PDSs in the nonabelian group $C_{p^{d}}\rtimes C_{(p^{d}-1)/6}\leqslant({\rm GF}(p^{d}),+)\rtimes({\rm GF}(p^{d})^{\times},\cdot)$.
@@
Let $\alpha$ be the generator for $({\rm GF}(p^{d})^{\times},\cdot)$, and let $H=\langle\alpha^{(p^{d}-1)/k(k-1)}\rangle$. Let $B$ be a $k$ subset of $({\rm GF}(p^{d})^{\times},\cdot)$, and set =$\mathscr{B}=\{Bg:g\in H\}$. Consider the set of nonzero differences $\delta B=\{b_{i}-b_{j}:b_{i},b_{j}\in B,b_{i}\neq b_{j}\}$ evaluated over the field ${\rm GF}(p^{d})$. If $|\delta B\cap Hg|$ is constant for each coset $Hg$ in $({\rm GF}(p^{d})^{\times},\cdot)$, then by [33, Theorem 2], $\mathscr{B}$ forms the set of base blocks (defined in 2.3) for an ${\mathrm{S}}(2,k,p^{d})$. By [33, Corollary 1], there is always some set $B$ such that $\delta B$ is evenly distributed over the cosets $Hg$ whenever $p^{d}>(\frac{1}{2}k(k-1))^{k(k-1)}$.
@@
When $p^{d}\leqslant(\frac{1}{2}k(k-1))^{k(k-1)}$, Buratti [5] provides sufficient conditions for block regular ${\mathrm{S}}(2,k,p^{d})$; we mention one result here, which implies the existence of the (183,70,29,25)(183,70,29,25)( 183 , 70 , 29 , 25 )-PDS we found in SmallGroup(183, 1) related to a block regular ${\mathrm{S}}(2,5,61)$.
@@
Let $p^{d}=20t+1$ be a prime power, let $2^{e}$ be the highest power of 222 in $t$, and suppose
@@
Let $\alpha$ generate the multiplicative group $({\rm GF}(p^{d})^{\times},\cdot)$. By [5, Theorem 7], there always exists a block $B$ whose differences form a system of distinct representatives for the cosets of the group $H=\langle\alpha^{\frac{p-1}{20}}\rangle$. By [33, Theorem 2], the set =$\mathscr{B}=\{Bg:g\in H\}$ forms the set of base blocks for a block regular ${\mathrm{S}}(2,5,p)$. By the discussion in Subsection 2.3, this gives a $(p^{d}(p^{d}-1)/20,5(p^{d}-5)/4,(p^{d}-1)/4+14,25)$-PDS. ∎
@@
[…]
@@
[…]
@@ 4.4. Non-modular techniques
We will prove that there is no reversible (144,66,30)(144,66,30)( 144 , 66 , 30 )-DS $D$ in SmallGroup(144, 68). In this case, $\sqrt{\Delta}=12$, where $12^{2}=144$. Therefore, Theorem 3.8, Theorem 3.11, Theorem 3.13, and Algorithm CCI are all inapplicable.
Examining the character table for $G$ (which can be generated, for example, using GAP[15]), there are six conjugacy classes of noncentral elements of order 333. Each of these conjugacy classes has size 161616, so each element in one of these classes has a centralizer of order 999. Thus, for any noncentral element $x$ of order 333 in $G$, we have
@@
Moreover, if $\mathcal{L}$ denotes the set of linear characters for $G$, then $\chi\notin\mathcal{L}$ implies $\chi(x)=0$. Note that $|\mathcal{L}|=9$, and each nonprincipal character has order 333; other than the principal character, there are four conjugate pairs $\{\xi_{i},\overline{\xi_{i}}\},$ $1\leqslant i\leqslant 4$. Extracting the principal character from this sum and noting that $D$ is reversible (so $\overline{\chi(D)}=\chi(D)$ by Lemma 3.4(i)), we have
@@
Since each $\xi_{i}(D)$ is ±6$\pm$ 6± 6 by Lemma 3.2, there are only $2^{4}=16$ possibilities in total. Checking which have $\Phi(x)$ divisible by 999, there are only 555 possibilities that work. (In each case, $\Phi(x)=36$ or $\Phi(x)=72$.)
Of the 151515 remaining nonlinear characters, there is one that equals its conjugate and seven conjugate pairs. Let $\chi$ be a nonlinear irreducible character. Note that $\theta_{1}=6$ and $\theta_{2}=-6$. By Lemma 3.2, if $\mathfrak{F}$ is the representation affording $\chi$, then $\chi(D)=6(\dim V_{\mathfrak{F}}(6)-\dim V_{\mathfrak{F}}(-6))$, where $-\chi(1)\leqslant(\dim V_{\mathfrak{F}}(6)-\dim V_{\mathfrak{F}}(-6))\leqslant\chi(1)$. So, we may conclude that $\chi(D)\in\{-18,-6,6,18\}$.
@@
Since either $1\in D$ or $1\notin D$, we have $\Phi(1)\in\{0,144\}$. Of the $5\cdot 4^{8}$ possibilities for $\Phi$, a GAP [15] calculation shows that only 158201582015820 of these possibilites have $\Phi(1)\in\{0,144\}$.
Now, $Z(G)=\langle z\rangle$ is cyclic of order 333. Again, since $D$ is reversible, either $\{z,z^{-1}\}\subseteq D$ or $\{z,z^{-1}\}\cap D=\varnothing$. Thus, either
@@
None of the remaining possibilities for $\Phi$ satisfy either of these constraints (in fact, all the values of these class functions are $6\pmod{12}$ when evaluated at $z$), and so no such $\Phi$ can exist. Therefore, there is no reversible (144,66,30)(144,66,30)( 144 , 66 , 30 )-DS (and hence no regular (144,66,30,30)(144,66,30,30)( 144 , 66 , 30 , 30 )-PDS) in $G$.
@@ Acknowledgements
Appendix A Parameters Ruled using Theorem 3.13
The following tables include all parameters $(v,k,\lambda,\mu)$ with $k<(v-1)/2$ listed in Brouwer’s tables [3] that can be ruled out using Theorem 3.13 on either the original parameters or those of its complement, excluding those parameters with $v\in\{512,768,1024\}$, in which cases the computations are too expensive due to the number of groups on those orders. Note, we do not list both a parameter set and its complement.
@@
The first set of tables tables consist of all parameters $(v,k,\lambda,\mu)$ with $k<(v-1)/2$ listed in Brouwer’s tables [3] satisfying the requirement that $\Delta$ is divisible by some prime $p$ not dividing $v$, for which we have shown no PDS exists through direct computation using Algorithm CCI on either the original parameters or those of the complement. We have only included all parameters with $v<507$, for otherwise the computations become too expensive. Once again, we do not include complements of listed parameters.
@@
We have additionally computed all parameters $(v,k,\lambda,\mu)$ listed in Brouwer’s tables with $k<(v-1)/2$ and $507\leqslant v\leqslant 1300$ such that $\gcd(v,\Delta)=1$. Of those, the following parameter sets are ruled out by our computations using Algorithm CCI.
@@
Finally, we mention again here that the parameter sets (111,30,5,9)(111,30,5,9)( 111 , 30 , 5 , 9 ) and (96,35,10,14)(96,35,10,14)( 96 , 35 , 10 , 14 ) have been ruled out using a combination of Algorithm CCI and Gurobi [19] (Remark 4.8).
@@ Appendix C Parameter sets with a valid output from Algorithm CCI
We now list all parameter sets $(v,k,\lambda,\mu)$ with $k<(v-1)/2$ listed in Brouwer’s tables [3] satisfying the condition that either $\Delta$ has some distinct prime factor with $v$ and $v\leqslant 507$ or $\gcd(v,\Delta)=1$ and $v\leqslant 1300$, such that there does exist a valid $\Phi$ function, or such that this function was not computable with the available computational resources. We note once again that we exclude parameter sets with $v\in\{512,768,1024\}$, as well as the parameter set (507,176,70,56)(507,176,70,56)( 507 , 176 , 70 , 56 ), for these computations are too expensive. As mentioned, we do not include complements of listed parameters – these may be easily generated from what is provided. We begin with the table of all parameters in groups which were too computationally expensive to compute.
@@
[…]

### v2

@@ 4.1. Modular computation of conjugacy class intersection sizes for PDSs
$[0,9,1,9,1,1,1,1,1,1,1,1,1,1,1]$
as a potentially valid list of conjugacy class intersection sizes for a $(111,30,5,9)$-PDS in SmallGroup(111, 1). Note that a subset of this group with these intersection sizes will evaluate on irreducible characters to values admissible by Lemma 3.2. However, by turning the conjugacy class and PDS constraints into a linear program, we are able to rule out the existence of such a PDS using the linear optimization software Gurobi [19], and hence there does not exist a $(111,30,5,9)$-PDS.
@@
[0,4,14,1,1,4,4,1,2,4],[0,4,14,1,1,4,4,2,1,4][0,4,14,1,1,4,4,1,2,4],[0,4,14,1,1,4,4,2,1,4]
as potentially valid lists of conjugacy class intersections for a $(96,35,10,14)$-PDS in SmallGroup(96, 64) and SmallGroup(96, 227), respectively. These were also ruled out with the use of Gurobi [19], and, since these are the only valid lists of conjugacy class intersection sizes for these parameters, there does not exist a $(96,35,10,14)$-PDS, either. The authors would be quite interested in a proof of the nonexistence of these PDSs that does not use the brute force of linear programming.
@@ 4.3. PDSs constructed using Algorithm CCI
For a number of groups, we used this method to calculate all possible sizes of intersections of the conjugacy classes of the group with a possible PDS inside this group. We then ran a restricted search for PDSs satisfying the class intersection restrictions using a modified Random Restart Hill Climb search described in [2]. Using the technique from [2], we were able to construct nineteen distinct PDSs in nonabelian groups, many of which were already explicitly recognized, such as PDSs arising from triangular graphs. We found three parameter sets, $(111,44,19,16)$, $(305,76,27,16)$, and $(981,140,43,16)$, which correspond to an infinite family of groups acting regularly on Steiner systems ${\mathrm{S}}(2,4,p^{d})$ for $p$ prime. Five others that we were able to construct computationally – in SmallGroup(57, 1), SmallGroup(155, 1), SmallGroup(301, 1), SmallGroup(737, 1), and SmallGroup(1027, 1) – have parameters corresponding to an infinite family of PDSs corresponding to Steiner triple systems. The infinite families on ${\mathrm{S}}(2,4,p)$ and ${\mathrm{S}}(2,3,p^{d})$ correspond to block-regular Steiner $2$-design constructions given by Clapham in [6] and Wilson in [33]. Finally, we found a $(183,70,29,25)$-PDS in SmallGroup(183, 1) which is related to a ${\mathrm{S}}(2,5,61)$ Steiner system and, as it turns out, is also part of an infinite family. To the authors knowledge, the PDSs arising from [33] have not previously been recognized in the literature, and we are only aware of one paper [28] recognizing the constructions given by Clapham [6] as giving rise to PDSs.
@@
Let $p$ be a prime such that $p^{d}>9$ and $p^{d}\equiv 7\pmod{12}$. Then, there exists a $(p^{d}(p^{d}-1)/6,3(p^{d}-3)/2,(p^{d}+3)/2,9)$-PDS in the nonabelian group $C_{p^{d}}\rtimes C_{(p^{d}-1)/6}\leqslant({\rm GF}(p^{d}),+)\rtimes({\rm GF}(p^{d})^{\times},\cdot)$.
@@
Let $\alpha$ be the generator for $({\rm GF}(p^{d})^{\times},\cdot)$, and let $H=\langle\alpha^{(p^{d}-1)/k(k-1)}\rangle$. Let $B$ be a $k$ subset of $({\rm GF}(p^{d})^{\times},\cdot)$, and set $\mathscr{B}=\{Bg:g\in H\}$. Consider the set of nonzero differences $\delta B=\{b_{i}-b_{j}:b_{i},b_{j}\in B,b_{i}\neq b_{j}\}$ evaluated over the field ${\rm GF}(p^{d})$. If $|\delta B\cap Hg|$ is constant for each coset $Hg$ in $({\rm GF}(p^{d})^{\times},\cdot)$, then by [33, Theorem 2], $\mathscr{B}$ forms the set of base blocks (defined in 2.3) for an ${\mathrm{S}}(2,k,p^{d})$. By [33, Corollary 1], there is always some set $B$ such that $\delta B$ is evenly distributed over the cosets $Hg$ whenever $p^{d}>(\frac{1}{2}k(k-1))^{k(k-1)}$.
@@
When $p^{d}\leqslant(\frac{1}{2}k(k-1))^{k(k-1)}$, Buratti [5] provides sufficient conditions for block regular ${\mathrm{S}}(2,k,p^{d})$; we mention one result here, which implies the existence of the $(183,70,29,25)$-PDS we found in SmallGroup(183, 1) related to a block regular ${\mathrm{S}}(2,5,61)$.
@@
Let $p^{d}=20t+1$ be a prime power, let $2^{e}$ be the highest power of $2$ in $t$, and suppose
@@
Let $\alpha$ generate the multiplicative group $({\rm GF}(p^{d})^{\times},\cdot)$. By [5, Theorem 7], there always exists a block $B$ whose differences form a system of distinct representatives for the cosets of the group $H=\langle\alpha^{\frac{p-1}{20}}\rangle$. By [33, Theorem 2], the set $\mathscr{B}=\{Bg:g\in H\}$ forms the set of base blocks for a block regular ${\mathrm{S}}(2,5,p)$. By the discussion in Subsection 2.3, this gives a $(p^{d}(p^{d}-1)/20,5(p^{d}-5)/4,(p^{d}-1)/4+14,25)$-PDS. ∎
@@
If $p\equiv 13\pmod{2k(k-1)}$ and $p\neq 13$, then there exists a (p(p1)/12,4(p4)/3,(p+20)/3,16)(p(p-1)/12,4(p-4)/3,(p+20)/3,16)-PDS inside the nonabelian group $C_{p}\rtimes C_{(p-1)/12}\leqslant({\rm GF}(p),+)\rtimes({\rm GF}(p)^{\times},\cdot)$.
@@
Let $\alpha$ generate the multiplicative group $({\rm GF}(p)^{\times},\cdot)$. By [14, Theorem 3.9], there always exists a block $B$ whose differences form a system of distinct representatives for the cosets of the group $H=\langle\alpha^{\frac{p-1}{12}}\rangle$. By [33, Theorem 2], the set $\mathscr{B}=\{Bg:g\in H\}$ forms the set of base blocks for a block regular ${\mathrm{S}}(2,4,p)$. By the discussion in Subsection 2.3, this gives a (p(p1)/12,4(p4)/3,(p+20)/3,16)(p(p-1)/12,4(p-4)/3,(p+20)/3,16)-PDS. ∎
@@ 4.4. Non-modular techniques
We will prove that there is no reversible $(144,66,30)$-DS $D$ in SmallGroup(144, 68). In this case, $\sqrt{\Delta}=12$, where $12^{2}=144$. Therefore, Theorem 3.8, Theorem 3.12, Theorem 3.14, and Algorithm CCI are all inapplicable.
Examining the character table for $G$ (which can be generated, for example, using GAP[15]), there are six conjugacy classes of noncentral elements of order $3$. Each of these conjugacy classes has size $16$, so each element in one of these classes has a centralizer of order $9$. Thus, for any noncentral element $x$ of order $3$ in $G$, we have
@@
Moreover, if $\mathcal{L}$ denotes the set of linear characters for $G$, then $\chi\notin\mathcal{L}$ implies $\chi(x)=0$. Note that $|\mathcal{L}|=9$, and each nonprincipal character has order $3$; other than the principal character, there are four conjugate pairs $\{\xi_{i},\overline{\xi_{i}}\},$ $1\leqslant i\leqslant 4$. Extracting the principal character from this sum and noting that $D$ is reversible (so $\overline{\chi(D)}=\chi(D)$ by Lemma 3.4(i)), we have
@@
Since each $\xi_{i}(D)$ is $\pm 6$ by Lemma 3.2, there are only $2^{4}=16$ possibilities in total. Checking which have $\Phi(x)$ divisible by $9$, there are only $5$ possibilities that work. (In each case, $\Phi(x)=36$ or $\Phi(x)=72$.)
Of the $15$ remaining nonlinear characters, there is one that equals its conjugate and seven conjugate pairs. Let $\chi$ be a nonlinear irreducible character. Note that $\theta_{1}=6$ and $\theta_{2}=-6$. By Lemma 3.2, if $\mathfrak{F}$ is the representation affording $\chi$, then $\chi(D)=6(\dim V_{\mathfrak{F}}(6)-\dim V_{\mathfrak{F}}(-6))$, where $\chi(1)\leqslant(\dim V_{\mathfrak{F}}(6)-\dim V_{\mathfrak{F}}(-6))\leqslant\chi(1)$. So, we may conclude that $\chi(D)\in\{-18,-6,6,18\}$.
@@
Since either $1\in D$ or $1\notin D$, we have $\Phi(1)\in\{0,144\}$. Of the $5\cdot 4^{8}$ possibilities for $\Phi$, a GAP [15] calculation shows that only $15820$ of these possibilites have $\Phi(1)\in\{0,144\}$.
Now, $Z(G)=\langle z\rangle$ is cyclic of order $3$. Again, since $D$ is reversible, either $\{z,z^{-1}\}\subseteq D$ or $\{z,z^{-1}\}\cap D=\varnothing$. Thus, either
@@
None of the remaining possibilities for $\Phi$ satisfy either of these constraints (in fact, all the values of these class functions are $6\pmod{12}$ when evaluated at $z$), and so no such $\Phi$ can exist. Therefore, there is no reversible $(144,66,30)$-DS (and hence no regular $(144,66,30,30)$-PDS) in $G$.
@@ Acknowledgements
Appendix A Parameters already ruled out by [32]
The following tables include all parameters $(v,k,\lambda,\mu)$ with k<(v1)/2k<(v-1)/2 listed in Brouwer’s tables [3] that can be ruled out (entirely, in any group) using either Theorem 3.14 or [32, Corollaries 5.2, 5.3] on either the original parameters or those of its complement, excluding those parameters with $v\in\{512,768,1024\}$, in which cases the computations are too expensive due to the number of groups on those orders. Note, we do not list both a parameter set and its complement.
@@ Appendix A Parameters already ruled out by [32]
The first set of tables tables consist of all parameters $(v,k,\lambda,\mu)$ with k<(v1)/2k<(v-1)/2 listed in Brouwer’s tables [3] satisfying the requirement that $\Delta$ is divisible by some prime $p$ not dividing $v$, for which we have shown no PDS exists in any group through direct computation using Algorithm CCI on either the original parameters or those of the complement. We have only included all parameters with $v<507$, for otherwise the computations become too expensive. Once again, we do not include complements of listed parameters.
@@
We have additionally computed all parameters $(v,k,\lambda,\mu)$ listed in Brouwer’s tables with k<(v1)/2k<(v-1)/2 and $507\leqslant v\leqslant 1300$ such that $\gcd(v,\Delta)=1$. Of those, the following parameter sets are ruled out by our computations using Algorithm CCI.
@@
Finally, we mention again here that the parameter sets $(111,30,5,9)$ and $(96,35,10,14)$ have been ruled out using a combination of Algorithm CCI and Gurobi [19] (Remark 4.8).
@@ Appendix C Parameter sets with a valid output from Algorithm CCI
We now list all parameter sets $(v,k,\lambda,\mu)$ with k<(v1)/2k<(v-1)/2 listed in Brouwer’s tables [3] satisfying the condition that either $\Delta$ has some distinct prime factor with $v$ and $v\leqslant 507$ or $\gcd(v,\Delta)=1$ and $v\leqslant 1300$, such that there does exist a valid $\Phi$ function, or such that this function was not computable with the available computational resources. We note once again that we exclude parameter sets with $v\in\{512,768,1024\}$, as well as the parameter set $(507,176,70,56)$, for these computations are too expensive. As mentioned, we do not include complements of listed parameters – these may be easily generated from what is provided. We begin with the table of all parameters in groups which were too computationally expensive to compute.
@@
We now provide all successful computations. We have organized the table as follows. First by the parameter set $(v,k,\lambda,\mu)$, then by the column labeled ID, which provides the GAP [15] group ID for which a valid $\Phi$ function has been constructed, then by the column labeled ?, which describes if a PDS for that group on that parameter set has been constructed (in which case a reference is given to its construction, either in this paper or elsewhere) or has not been either constructed or ruled out (?), in which the existence is still an open question. Finally, we provide the conjugacy class intersections of the possible PDS inside of the group. The intersections are given in the form of a list $a_{1},\dots,a_{n}$], such that, if $G$ is the group, then the $i$th entry of $a_{1},\dots,a_{n}$] corresponds to the intersection of the $i$th conjugacy class in ConjugacyClasses(CharacterTable($G$)) with the possible PDS. (Note that ConjugacyClasses($G$) and ConjugacyClasses(CharacterTable($G$)) may have different conjugacy class orderings.) We have listed all conjugacy class intersections up to the automorphisms of $G$. That is, if a conjugacy class intersection is omitted for a given group, then an automorphically equivalent intersection is listed. We also provide a full .txt file containing all of these parameters in this GitHub repository.
We also mention that the parameter set $(243,112,46,56)$ has over 400 valid, automorphically distinct intersections. To avoid cluttering the paper, these have been placed in a separate file at the same GitHub repository located here.